OpenStudy (anonymous):

simplest form of the expression 3sqrt750+3sqrt2058-3sqrt48 can someone explain

OpenStudy (shamim):

\[3\sqrt{750}+3\sqrt{2058}-3\sqrt{48}\]

OpenStudy (anonymous):

yes thats what it looks like

OpenStudy (shamim):

ok

OpenStudy (shamim):

so u r now able to solve it. is not it

OpenStudy (anonymous):

no thats the equation possible answers are a4 3sqrt6 17.5 3sqrt6 33 3sqrt6 10 3sqrt6 I dont know how to solve it

OpenStudy (shamim):

ok

OpenStudy (shamim):

i m trying it

OpenStudy (shamim):

do u see ur question once again

OpenStudy (anonymous):

yes

OpenStudy (shamim):

do u think ur question is ok?

OpenStudy (anonymous):

what do you mean

OpenStudy (shamim):

i m not getting any result from ur question which will match with 1 of ur 4 options

OpenStudy (anonymous):

sorry it wants to know what is the simpliest form of the expression

OpenStudy (shamim):

understood

OpenStudy (shamim):

i can tell u the way to make it simplest

OpenStudy (shamim):

\[=3\sqrt{30 \times 25}+3\sqrt{6 \times 343}-3\sqrt{3 \times 16}\]

OpenStudy (shamim):

\[=3 \times 5\sqrt{30}+3\sqrt{6 \times 343}-3 \times 4\sqrt{3}\]

OpenStudy (shamim):

m i right

OpenStudy (anonymous):

thats not an answer on my sheet

OpenStudy (shamim):

ya

OpenStudy (shamim):

i said it b4

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Doesn't help!!

OpenStudy (shamim):

i think ur question may b wrong

OpenStudy (anonymous):

the answer was 10 3sqrt6, dont ask me im lost on this stuff, but thanks for trying

OpenStudy (anonymous):

Thank you @Lilman95!